# in Post-Training and Test-Time Scaling of Large Language Models

### v1

@@ Abstract
Recent developments in Large Language Models (LLMs) have shifted from pre-training scaling to post-training and test-time scaling. Across these developments, a key unified paradigm has arisen: Learning from Rewards, where reward signals act as the guiding stars to steer LLM behavior. It has underpinned a wide range of prevalent techniques, such as reinforcement learning (in RLHF, DPO, and GRPO), reward-guided decoding, and post-hoc correction. Crucially, this paradigm enables the transition from passive learning from static data to active learning from dynamic feedback. This endows LLMs with aligned preferences and deep reasoning capabilities. In this survey, we present a comprehensive overview of the paradigm of learning from rewards. We categorize and analyze the strategies under this paradigm across training, inference, and post-inference stages. We further discuss the benchmarks for reward models and the primary applications. Finally we highlight the challenges and future directions. We maintain a paper collection at GitHub repository.
[…]
citnum
Sailing AI by the Stars: A Survey of Learning from Rewards
in Post-Training and Test-Time Scaling of Large Language Models
Xiaobao Wu Nanyang Technological University xiaobao.wu@ntu.edu.sg
@@ 1 Introduction
Recent years have witnessed the rapid advancement of Large Language Models (LLMs), such as ChatGPT OpenAI (2023), Claude Anthropic (2025), and Llama Meta (2023, 2024). These models are initially empowered by pre-training scaling Kaplan et al. (2020), which trains LLMs on massive corpora through next-token prediction. While this approach enables broad linguistic and knowledge representations, it suffers from several fundamental limitations: misalignment with human values Bai et al. (2022b); Zhang et al. (2023b); Deshpande et al. (2023), difficulty in adapting to various task objectives Lyu et al. (2023); Wang et al. (2023a), and deficiencies in deep reasoning Mirzadeh et al. (2024); Wu et al. (2024h). As a result, these limitations confine pre-trained models to surface-level tasks, falling short of the long-term goal of robust and general AI. To address these limitations, recent efforts have turned to post-training and test-time scaling, which seek to further refine LLMs after pre-training.
@@
Across the post-training and test-time scaling, a critical unified paradigm has emerged as illustrated in Figure 1: Learning from Rewards, which leverages reward signals to guide model behavior through diverse learning strategies. For post-training scaling, this paradigm has underpinned several key techniques, including preference alignment through Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022) or AI Feedback (RLAIF, Bai et al., 2022b) with scalar rewards and PPO Schulman et al. (2017), and Direct Preference Optimization (DPO, Rafailov et al., 2023) with implicit rewards. For test-time scaling, this paradigm supports eliciting long Chain-of-Thoughts reasoning via GRPO Shao et al. (2024) with rule-based rewards in DeepSeek-R1 DeepSeek-AI et al. (2025), generate-then-rank (Best-of-N) Cobbe et al. (2021); Lightman et al. (2023), reward-guided decoding Deng and Raffel (2023); Khanov et al. (2024), and post-hoc correction Akyurek et al. (2023); Madaan et al. (2023). Through these techniques, this paradigm enables LLMs to learn actively from dynamic feedback, in contrast to learning passively from static data. As such, this endows LLMs with aligned preferences and deep reasoning and planning abilities, leading to more intelligent agents. In consequence, this paradigm has inspired many applications, such as mathematical reasoning Wang et al. (2023b); DeepSeek-AI (2025), code generation Zhu et al. (2024); Zhou et al. (2025a), multimodality Liu et al. (2025h), agents Xia et al. (2025); OpenAI (2025), and embodied AI Zhang et al. (2025d); Zhao et al. (2025a).
@@
Due to this growing prevalence, we in this paper comprehensively review the learning from rewards for LLMs. We first introduce a taxonomy that categorizes existing works with a unified conceptual framework regarding reward model design and learning strategies. Then we review representative techniques across three main stages: training with rewards, inference with rewards, and post-inference with rewards. We additionally summarize recent reward model benchmarks and finally conclude by outlining key challenges and promising directions for future research.
@@ 2 A Taxonomy of Learning from Rewards for LLMs
We first introduce a unified conceptual framework that captures the key components and interactions to understand learning from rewards systemically. Building upon this framework, we then categorize the primary dimensions along which existing methods vary: (i) What is the source of rewards; (i) How to design the reward model; (i) When to learn from rewards; (i) How to learn from rewards.
@@ 2.1 A Unified Conceptual Framework
Figure 3: Reward Model (RM) design dimensions: (a) Model Architecture (Model-based and Model-free); (b) Reward Format (Scalar, Critique, and Implicit); (c) Scoring Pattern (Pointwise and Pairwise); (d) Reward Granularity (Outcome and Process).
2.2 What is the Source of Rewards?
@@ 2.3 How to Design the Reward Model?
Designing the reward model is the central foundation of learning from rewards. As shown in Figure 3, we organize the design space into four key dimensions as follows.
  • Base Architecture. As shown in Figure 3(a), this refers to the base architecture of a reward model: (i) Model-based. A dedicated reward model is trained to evaluate outputs. Variants include scalar reward models (outputting numerical scores), generative reward models (producing critiques), and semi-scalar reward models (combining both). Modern rewards models generally adopt the Transformers framework Liu et al. (2024a). (i) Model-free. Reward signals are directly modeled from the source without an explicit trained model, such as DPO Rafailov et al. (2023) and rule-based rewards of GRPO DeepSeek-AI et al. (2025). In order to align with previous literature, we hereafter refer to the reward model as the model-based by default.
@@
  • Reward Format. As shown in Figure 3(b), this describes the specific format of reward signals: (i) Scalar rewards, numerical scores that quantify the quality of model outputs. They are the most commonly used format due to their simplicity and compatibility with learning strategies such as reinforcement learning. Their limitation lies in the sparsity and interpretability. (i) Critique rewards, natural language feedback that evaluates the quality of outputs Saunders et al. (2022); Kwon et al. (2023), such as “The score of this response is 3 out of 5”. They are more expressive and interpretable than scalar rewards, enabling finer-grained guidance, but they may require additional processing to be used in some learning algorithms. (i) Implicit rewardsare signals implicitly embedded in the source without explicit supervision, such as preference data in DPO Rafailov et al. (2023). This format simplifies the implementation but places more burden on the learning strategies to infer appropriate optimization signals.
@@
  • Scoring Pattern. As shown in Figure 3(c), this dimension determines how responses are scored: (i) Pointwise, assigning a score to each response independently; (i) Pairwise, comparing response pairs and selecting the preferred one. The pairwise scoring can be expressed as a scalar score indicating relative preference or a natural language critique such as “Response 1 is better than Response 2”.
  • Reward Granularity. As shown in Figure 3(d), we identify two kinds of reward granularity: (i) Outcome-level, evaluating the holistic quality of outputs. (i) Process-level, assigning rewards to the intermediate steps within the reasoning process of outputs.
2.4 When to Learn from Rewards?
@@
Learning from rewards can occur at different stages of the language model lifecycle:
@@
2.5 How to Learn from Rewards?
Various learning approaches have been developed to incorporate reward signals to steer model behavior. These approaches are commonly divided into two types: training-based and training-free.
  • Training-based Strategies. Training-based strategies adopt reward signals to optimize the language model parameters. The optimization mainly depends on Reinforcement Learning (RL) and Supervised Fine-Tuning (SFT). Representative examples include Proximal Policy Optimization (PPO, Schulman et al., 2017; Ouyang et al., 2022), Direct Preference Optimization (DPO, Rafailov et al., 2023; Meng et al., 2024), Group Relative Policy Optimization (GRPO, Shao et al., 2024), and Rejection-Sampling Fine-Tuning (RSFT, Nakano et al., 2021; Yuan et al., 2023a; Dong et al., 2023)
  • Training-free Strategies. Training-free strategies leverage reward signals to guide or refine model outputs without updating the language model parameters. They include generate-then-rank, such as Best-of-N Cobbe et al. (2021); Lightman et al. (2023), reward-guided decoding Deng and Raffel (2023); Khanov et al. (2024), and post-inference correction Shinn et al. (2023); Pan et al. (2023a). These methods offer a relatively lightweight mechanism to improve model outputs; some are highly compatible with different model architectures. They are particularly useful when model fine-tuning is infeasible or computationally expensive.
The above methods provide distinct mechanisms to learn the language model from reward signals. We will review them in detail in the following § 3, 4 and 5.
@@ 3 Training with Rewards
Fine-Grained RLHF Wu et al. (2023e); Fact-RLHF Sun et al. (2023);
@@
Critic-RM Yu et al. (2024b); MM-RLHF Zhang et al. (2025f) , leaf, fill=training-fill, draw=none, text width= ] ] [ Training with
@@
RLHF-V Yu et al. (2023b); UnifiedRM Wang et al. (2025d);
@@ 3.1 Training with Scalar Rewards
Various variants have been explored based on RLHF. Safe RLHF Dai et al. (2023) emphasizes safety-centric alignment. Fine-Grained RLHF Wu et al. (2023e) moves beyond holistic feedback to dense, segment-level fine-grained feedback regarding relevance, factual accuracy, and completeness.
Training with scalar rewards has also been extended to multimodal tasks. Fact-RLHF Sun et al. (2023) incorporates factuality-aware scalar rewards to reduce hallucinations of multimodal understanding. Others focus on multimodal generation tasks. HPS and its variant Wu et al. (2023d, c) train a binary classifier as the reward model to evaluate generated image quality. Lee et al. (2023b) trains a multimodal reward model based on human feedback and fine-tunes a diffusion model to improve image generation. ImageReward Xu et al. (2023) introduces a general-purpose reward model for text-to-image generation. It is trained on image preferences from human experts and then fine-tunes a diffusion model via a direct tuning algorithm. RAHF Liang et al. (2024) further enriches the reward signals with scalar scores, heatmaps, and keyword omissions to guide sample filtering for fine-tuning. LiFT Wang et al. (2024f) applies a similar idea to the text-to-video task.
@@ 3.2 Training with Critique Rewards
Furthermore, several methods adopt a hybrid structure—semi-scalar reward models that produce both scalar and critique rewards. Early attempts incorporate critiques into the training of a scalar reward model, such as CLoud Ankner et al. (2024), MATH-Minos Gao et al. (2024a), and Ye et al. (2024a). Critic-RM Yu et al. (2024b) is a reward model with two heads, one for generating critiques and another for predicting scalars. It is trained on self-generated critiques of model outputs through a multi-task objective. MM-RLHF Zhang et al. (2025f) extends it to the multimodal field. It trains a semi-scalar reward model on the image, video understanding, and safety and then fine-tunes a multimodal language model with the trained reward model.
@@ 3.3 Training with Implicit Rewards
Some studies have extended training with implicit rewards in reasoning and multimodal areas. In mathematical reasoning, Step-DPO Lai et al. (2024) and Full-Step-DPO Xu et al. (2025a) build step-level preference data for DPO training to improve their reasoning abilities. In multimodal areas, mDPO Wang et al. (2024a) proposes a multimodal extension of DPO. It introduces conditional and anchored preference objectives to better leverage visual inputs and preserve the likelihood of preferred responses in multimodal tasks. RLHF-V Yu et al. (2023b) fine-tunes a multimodal LLM through DPO with human preferences regarding trustworthiness. UnifiedRM Wang et al. (2025d) trains a unified multimodal reward model on human preferences for several key tasks, including image/video understanding and generation. Later, it builds preference data through the trained reward model for DPO training. Beyond language models, these techniques are applied to diffusion models due to their simplicity, such as Diffusion-DPO Wallace et al. (2024).
@@
    A list of extended studies focuses on the multimodal field. HA-DPO Zhao et al. (2023) evaluates model responses of image understanding with another trained model as a hallucination detector to produce preference data. With these preference data, it fine-tunes the multimodal language models via DPO to reduce hallucination. Many studies follow a similar way to this, such as POVID Zhou et al. (2024c), AMP Zhang et al. (2024d) and RLAIF-V Yu et al. (2024a), and BPO Pi et al. (2024).

### v2

# Xiaobao Wu

@@ Abstract
Recent developments in Large Language Models (LLMs) have shifted from pre-training scaling to post-training and test-time scaling. Across these developments, a key unified paradigm has arisen: Learning from Rewards, where reward signals act as the guiding stars to steer LLM behavior. It has underpinned a wide range of prevalent techniques, such as reinforcement learning (RLHF, RLAIF, DPO, and GRPO), reward-guided decoding, and post-hoc correction. Crucially, this paradigm enables the transition from passive learning from static data to active learning from dynamic feedback. This endows LLMs with aligned preferences and deep reasoning capabilities for diverse tasks. In this survey, we present a comprehensive overview of learning from rewards, from the perspective of reward models and learning strategies across training, inference, and post-inference stages. We further discuss the benchmarks for reward models and the primary applications. Finally we highlight the challenges and future directions.  [^1] We maintain a paper collection at https://github.com/bobxwu/learning-from-rewards-llm-papers.
Sailing by the Stars: A Survey on Reward Models and Learning Strategies for Learning from Rewards
Xiaobao Wu Nanyang Technological University xiaobao002@e.ntu.edu.sg
@@ 1 Introduction
Recent years have witnessed the rapid advancement of Large Language Models (LLMs), such as ChatGPT OpenAI (2023), Claude Anthropic (2025), and Llama Meta (2023, 2024). These models are initially empowered by pre-training scaling Kaplan et al. (2020), which trains LLMs on massive corpora through next-token prediction. While this approach enables broad linguistic and knowledge representations, it suffers from several fundamental limitations: misalignment with human values Bai et al. (2022b); Zhang et al. (2023b); Deshpande et al. (2023), difficulty in adapting to various task objectives Lyu et al. (2023); Wang et al. (2023a), and deficiencies in deep reasoning Mirzadeh et al. (2024); Wu et al. (2024b). As a result, these limitations confine pre-trained models to surface-level tasks, falling short of the long-term goal of robust and general AI. To address these limitations, recent efforts have turned to post-training and test-time scaling, which seek to further refine LLMs after pre-training.
@@
Across the post-training and test-time scaling, a critical unified paradigm has emerged as illustrated in Figure 1: Learning from Rewards, which leverages reward signals to guide model behavior through diverse learning strategies. For post-training scaling, this paradigm has underpinned several key techniques, including preference alignment through Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022) or AI Feedback (RLAIF, Bai et al., 2022b) with scalar rewards and PPO Schulman et al. (2017), and Direct Preference Optimization (DPO, Rafailov et al., 2023) with implicit rewards. For test-time scaling, this paradigm supports eliciting long Chain-of-Thoughts reasoning via GRPO Shao et al. (2024) with rule-based rewards in DeepSeek-R1 DeepSeek-AI et al. (2025), generate-then-rank (Best-of-N) Cobbe et al. (2021); Lightman et al. (2023), reward-guided decoding Deng and Raffel (2023); Khanov et al. (2024), and post-hoc correction Akyurek et al. (2023); Madaan et al. (2023). By these techniques, this paradigm enables LLMs to learn actively from dynamic feedback, in contrast to learning passively from static data. As such, this endows LLMs with aligned preferences and deep reasoning and planning abilities, leading to more intelligent agents. In consequence, this paradigm has inspired many applications, such as mathematical reasoning Wang et al. (2023b); DeepSeek-AI (2025), code generation Zhu et al. (2024); Zhou et al. (2025a), multimodality Liu et al. (2025h), agents Xia et al. (2025); OpenAI (2025), and embodied AI Zhang et al. (2025d); Zhao et al. (2025a).
@@
Due to this growing prevalence, we in this paper comprehensively review the learning from rewards for LLMs. We first introduce a taxonomy that categorizes existing works with a unified conceptual framework regarding reward models and learning strategies. Then we review representative techniques across three main stages of LLMs: training with rewards, inference with rewards, and post-inference with rewards. We additionally summarize recent reward model benchmarks and finally conclude by outlining key challenges and promising directions for future research.
@@ 2 A Taxonomy of Learning from Rewards for LLMs
We first introduce a unified conceptual framework that captures the key components and interactions to understand learning from rewards systemically. Building upon this framework, we categorize the primary dimensions along which existing methods vary: (i) Reward Source; (ii) Reward Model; (iii) Learning Stage; (iv) Learning Strategies. Each dimension reflects a distinct aspect of how reward signals are acquired, represented, and utilized in language models.
@@ 2.1 A Unified Conceptual Framework
Figure 3: Reward Model (RM) design dimensions: (a) Base Architecture (Model-based and Model-free); (b) Reward Format (Scalar, Critique, and Implicit); (c) Scoring Pattern (Pointwise and Pairwise); (d) Reward Granularity (Outcome and Process).
2.2 Reward Source
@@ 2.3 Reward Model
Reward models are the central foundation of learning from rewards. As shown in Figure 3, we organize the design space of reward model into four key dimensions: (i) Base Architecture; (ii) Reward Format; (iii) Scoring Pattern; (iv) Reward Granularity.
Base Architecture.
As shown in Figure 3(a), this refers to the base architecture of a reward model. Here we consider a broad view of reward models, including both model-based and model-free architectures.
  • Model-based Architecture. A dedicated reward model is trained to evaluate outputs. Common variants include (a) Scalar Reward Models. These models assign a scalar score to a candidate response, indicating its quality. Typically, they are built upon Transformer backbones (e.g., GPT or BERT variants) with a value head that outputs scalars. They are trained with preference data via pairwise ranking losses such as the Bradley-Terry loss Nakano et al. (2021); Ouyang et al. (2022); Liu et al. (2024a).
    (a) Generative Reward Models. These models generate natural language critiques as reward signals. They commonly follow LLM-as-a-Judge with general models Zheng et al. (2023) or training specialized models Li et al. (2023a); Cao et al. (2024); Ye et al. (2024b); McAleese et al. (2024). They have become more popular recently because they can leverage the deep reasoning capabilities of large reasoning models and provide finer-grained supervision Huang et al. (2025a); Guo et al. (2025a).
    (a) Semi-scalar Reward Models. These models combine scalars with critiques, offering both quantitative and qualitative assessment Yu et al. (2024b); Zhang et al. (2025g). Their architectures usually involve two heads, one for scalar rewards and another for critique rewards.
  • Model-free Architecture. Instead of an explicit reward model, model-free approaches derive reward signals directly from diverse feedback sources, such as preference data, tools, or knowledge. The resulting rewards can be scalar, critique, or implicit signals. For example, DPO Rafailov et al. (2023) circumvents the need to train a reward model by directly aligning the language model with preference data through fine-tuning. Similarly, GRPO DeepSeek-AI et al. (2025) adopts rule-based rewards from handcrafted constraints and task-specific heuristics.
Model-based and model-free approaches each present distinct trade-offs in reward specification and practical applicability. Model-based approaches provide flexible and generalizable reward evaluation. Once trained, reward models can be reused across tasks and enable iterative optimization. However, they require costly preference data, are prone to overfitting, and may introduce bias or reward hacking issues. Model-free methods avoid training a separate reward model, offering a simpler, sample-efficient, and usually more stable pipeline. However, they are typically task-specific, lack generalization, and offer limited flexibility for downstream reuse.
In order to align with previous literature, we hereafter refer to the reward model as the model-based by default.
Reward Format.
As shown in Figure 3(b), this describes the specific format of reward signals:
  • Scalar Rewards, numerical scores that quantify the quality of model outputs. They are the most commonly used format due to their simplicity and compatibility with learning strategies such as reinforcement learning. Their limitation lies in the sparsity and interpretability.
@@
  • Critique Rewards, natural language feedback that evaluates the quality of outputs Saunders et al. (2022); Kwon et al. (2023), such as “The score of this response is 3 out of 5”. They are more expressive and interpretable than scalar rewards, enabling finer-grained guidance, but they may require additional processing to be used in certain learning algorithms.
  • Implicit Rewards are signals implicitly embedded in the source without explicit supervision, such as preference data in DPO Rafailov et al. (2023); Meng et al. (2024). This format simplifies the implementation but places more burden on the learning strategies to infer appropriate optimization signals.
Scoring Pattern.
As shown in Figure 3(c), this dimension determines how responses are scored:
  • Pointwise Scoring assigns a score to each response independently. It is the most widely used scoring pattern in reward models.
@@
  • Pairwise Scoring compares response pairs and selecting the preferred one. The pairwise scoring can be expressed as a scalar score indicating relative preference or a natural language critique such as “Response 1 is better than Response 2”.
Reward Granularity.
As shown in Figure 3(d), we identify two kinds of reward granularity: reward granularity reflects the level of resolution at which feedback is provided:
  • Outcome Reward Models evaluate the holistic quality of outputs, treating it as a single unit.
  • Process Reward Models evaluate intermediate steps within the reasoning process of outputs, enabling fine-grained supervision during generation Lightman et al. (2023); Wang et al. (2023b).
2.4 Learning Stage
@@
Learning from rewards can occur at different stages of the language model lifecycle, including Training, Inference, and Post-Inference.
@@
2.5 Learning Strategy
Various learning strategies have been developed to incorporate reward signals to steer model behavior. These strategies are commonly divided into two types: Training-based and Training-free.
  • Training-based Strategies. Training-based strategies optimize the language model by converting reward signals into gradient-based updates. The optimization mainly depends on Reinforcement Learning (RL) where language models act as policy models, or Supervised Fine-Tuning (SFT). Representative examples include Proximal Policy Optimization (PPO, Schulman et al., 2017; Ouyang et al., 2022), Direct Preference Optimization (DPO, Rafailov et al., 2023; Meng et al., 2024), Group Relative Policy Optimization (GRPO, Shao et al., 2024), and Rejection-Sampling Fine-Tuning (RSFT, Nakano et al., 2021; Yuan et al., 2023a; Dong et al., 2023)
  • Training-free Strategies. Training-free strategies leverage reward signals to guide or refine model outputs without updating the language model parameters. They include generate-then-rank, such as Best-of-N Cobbe et al. (2021); Lightman et al. (2023), reward-guided decoding Deng and Raffel (2023); Khanov et al. (2024), and post-inference correction Shinn et al. (2023); Pan et al. (2023a). These methods provide a relatively lightweight mechanism for improving model outputs, and some are highly compatible with various model architectures. They are particularly useful when model fine-tuning is infeasible or computationally expensive.
The above presents a detailed taxonomy of learning from rewards for LLMs. We will review the representative studies across the three learning stages: training, inference, and post-inference with rewards in the following § 3, 4 and 5.
@@ 3 Training with Rewards
Fine-Grained RLHF Wu et al. (2023c); Fact-RLHF Sun et al. (2023);
@@
RRM Guo et al. (2025a); Think-J Huang et al. (2025a);
UnifiedReward-Think Wang et al. (2025d) , leaf, fill=training-fill, draw=none, text width= ] ] [ Training with
Hybrid Rewards
§ 3.2, fill=training-fill, draw=training-border [ Critic-RM Yu et al. (2024b); MM-RLHF Zhang et al. (2025g);
Critique-GRPO Zhang et al. (2025e) , leaf, fill=training-fill, draw=none, text width= ] ] [ Training with
@@
RLHF-V Yu et al. (2023b); UnifiedRM Wang et al. (2025e);
@@ 3.1 Training with Scalar Rewards
Various variants have been explored based on RLHF. Safe RLHF Dai et al. (2023) emphasizes safety-centric alignment. Fine-Grained RLHF Wu et al. (2023c) moves beyond holistic feedback to dense, segment-level fine-grained feedback regarding relevance, factual accuracy, and completeness.
Training with scalar rewards has also been extended to multimodal tasks. Fact-RLHF Sun et al. (2023) incorporates factuality-aware scalar rewards to reduce hallucinations of multimodal understanding. Others focus on multimodal generation tasks. HPS and its variant Wu et al. (2023b, a) train a binary classifier as the reward model to evaluate generated image quality. Lee et al. (2023b) trains a multimodal reward model based on human feedback and fine-tunes a diffusion model to improve image generation. ImageReward Xu et al. (2023) introduces a general-purpose reward model for text-to-image generation. It is trained on image preferences from human experts and then fine-tunes a diffusion model via a direct tuning algorithm. RAHF Liang et al. (2024) further enriches the reward signals with scalar scores, heatmaps, and keyword omissions to guide sample filtering for fine-tuning. LiFT Wang et al. (2024f) applies a similar idea to the text-to-video task.
@@ 3.2 Training with Critique Rewards
3.3 Training with Hybrid Rewards
Furthermore, several methods adopt hybrid rewards (or semi-scalar rewards) that include both scalars and critiques. Early attempts incorporate critiques into the training of a scalar reward model, such as CLoud Ankner et al. (2024), MATH-Minos Gao et al. (2024a), and Ye et al. (2024a). Critic-RM Yu et al. (2024b) is a reward model with two heads, one for generating critiques and another for predicting scalars. It is trained on self-generated critiques of model outputs through a multi-task objective. Critique-GRPO Zhang et al. (2025e) applies both scalar and critique rewards for reinforcement learning. MM-RLHF Zhang et al. (2025g) extends it to the multimodal field. It trains a semi-scalar reward model on the image, video understanding, and safety and then fine-tunes a multimodal language model with the trained reward model.
@@ 3.4 Training with Implicit Rewards
Some studies have extended training with implicit rewards in reasoning and multimodal areas. In mathematical reasoning, Step-DPO Lai et al. (2024) and Full-Step-DPO Xu et al. (2025a) build step-level preference data for DPO training to improve their reasoning abilities. In multimodal areas, mDPO Wang et al. (2024a) proposes a multimodal extension of DPO. It introduces conditional and anchored preference objectives to better leverage visual inputs and preserve the likelihood of preferred responses in multimodal tasks. RLHF-V Yu et al. (2023b) fine-tunes a multimodal LLM through DPO with human preferences regarding trustworthiness. UnifiedRM Wang et al. (2025e) trains a unified multimodal reward model on human preferences for several key tasks, including image/video understanding and generation. Later, it builds preference data through the trained reward model for DPO training. Beyond language models, these techniques are applied to diffusion models due to their simplicity, such as Diffusion-DPO Wallace et al. (2024).
@@
    A list of extended studies focuses on the multimodal field. HA-DPO Zhao et al. (2023) evaluates model responses of image understanding with another trained model as a hallucination detector to produce preference data. With these preference data, it fine-tunes the multimodal language models via DPO to reduce hallucination. Many studies follow a similar way to this, such as POVID Zhou et al. (2024b), AMP Zhang et al. (2024d) and RLAIF-V Yu et al. (2024a), and BPO Pi et al. (2024).